OpenStudy (anonymous):

HELP ME PLEASE@$!$@!@!^&@!@#$#$@@!^&!@^!@$#@ Four polynomial functions are shown below: f(x) = x^3 + 3x^2 - 4x - 12 g(x) = x^3 - 3x - 2 p(x) = x^2 + 3x + 2 t(x) = x^2 + x - 2 Which of the following functions has 1 and -2 as its zeros? a.) f(x) b.)g(x) c.) p(x) d.) t(x)

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@zimmah

OpenStudy (anonymous):

@annas

OpenStudy (anonymous):

@Microrobot

OpenStudy (anonymous):

@ksaimouli

OpenStudy (ksaimouli):

graph them see which one touches the x-axis

OpenStudy (ksaimouli):

so b

OpenStudy (luigi0210):

Or work backwards: So we know \[\LARGE x=1, -2\] \[\LARGE (x=1)(x=-2)\] \[\LARGE (x-1)(x+2)=0\] Now distribute. btw, 2 roots=x^2 not x^3

OpenStudy (anonymous):

THANK'S

OpenStudy (luigi0210):

So not "B" but if you wanna take his word for it, go for it.